# Simplified vine copula models: state of science and affairs

### v2

@@ 1 Introduction
The simplifying assumption has become an oddly polarizing issue within our community, although the tension is not reflected in the published scientific record. Papers highlighting its limitations are dismissed in peer review for relying on ‘unrealistic toy examples.’ On the other side, methodological works on simplified models face rejections with the recommendation to abandon research involving the simplifying assumption altogether because it might be the ‘next formula that kills Wall Street’ (referring to the famous article by Salmon,, 2012, on the Gaussian copula). To my knowledge, neither of these positions has been voiced, let alone justified in the public scientific record. This state of affairs is quite unfortunate, as it hinders scientific progress.
The main purpose of this note is to bring the debate into the open scientific discourse and to offer my own, hopefully uncontroversial, perspective on the matter. At the very least, it should provide a good basis for a more constructive discourse going forward. The key points are as follows:
@@ 3.1.3 Extremal dependence
It appears that the simplifying assumption is quite natural in the context of extremal dependence models. This should not be misinterpreted as a general endorsement of simplified vine copulas for estimating extreme event probabilities. Tail copulas are objects different from copulas and have to be fitted only on the most extreme subset of observations. A simplified vine model fit to all observations can still be off, because the estimated parameters may underrepresent the dependence strength in tail.
@@ 3.2 Empirical evidence for and against the simplifying assumption
The aforementioned works provide anecdotal evidence that whether and how severely the simplifying assumption is violated in practice depends on the data at hand. From a philosophical point of view, there isn’t much reason to believe it is exactly true—apart from the special case of conditional independence. Any consistent test will detect the smallest deviation if the sample size is large enough, and the above references all use at least several hundred observations. The same can be said about any parametric model for the dependence and related goodness-of-fit tests. That does not mean that simplified or parametric models are inherently bad. What is important is that we understand the limitations of our models and the consequences of their violations.
@@ 3.3 Issues with interpreting simplified models
Figure 4: Contour plots of estimated pair-copulas in a simplified vine fit to non-simplified data (left: strong violation, right: moderate violation). While $X_{2}$ and $X_{3}$ are conditionally independent given $X_{1},X_{4}$, the partial vine copula induces spurious positive dependence.
@@
  • Specify conditional copulas in the second tree with $c_{2,3|1}(\cdot\mid u)=c_{3,4|1}(\cdot\mid u)$ as a Gumbel copula with Kendall’s $\tau$ equal to $0.9(2u-1)$ (strong violation) or $0.4(2u-1)$ (moderate violation), rotating the copula by 90 degrees if $\tau<0$.
@@
We now simulate 10 0001000010\,00010 000 observations from the model and fit a nonparametric simplified vine copula using the rvinecopulib package (Nagler and Vatter,, 2024). The fitted model on the left shows strong positive dependence in the pair $(X_{2},X_{3})$ given $(X_{1},X_{4})$, even though they are conditionally independent. Also for the moderate violation (right plot), there is weak positive correlation with a Kendall’s $\tau$ of approximately 0.08, which passes usual statistical significance thresholds. Concluding that the variables are conditionally dependent would clearly be a mistake. For the moderate violation, the pair-copulas in the second tree further seem close to independence, because they average out positive and negative dependencies. Concluding that the variables are conditionally independent is also a mistake.
@@ 3.4 The approximation error of simplified models
The approximation error was analyzed formally in a seminal paper by Mroz et al., (2021). Their most surprising result is that the collection of simplified vine copulas is dense in the set of all copulas with respect to the uniform metric $d_{\infty}(C,D)=\sup_{\bm{u}\in[0,1]^{d}}|C(\bm{u})-D(\bm{u})|$. This means that the approximation error can be made arbitrarily small by choosing the right simplified model. We must not misinterpret this result as saying that the simplifying assumption is inconsequential in practice. The proof of the result is based on the fact that all empirical copulas can be represented as simplified vine copulas, with independence pair-copulas everywhere but for the first tree. The same argument would suggest that Markov tree distributions are dense in the set of all distributions. However, assuming universal conditional independence is clearly restrictive in a practical sense. And indeed, this all breaks down when more statistically meaningful metrics are considered: Mroz et al., (2021) show that simplified vine copulas are nowhere dense with respect to the topology induced by weak convergence, the total variation metric, and the Kullback-Leibler divergence.
@@
It would be an equally big mistake to dismiss all models making the simplifying assumption based on these findings. The results are missing important context about alternative models. For example, I am not aware of similar results providing the approximation error of (hierarchical) Archimedean models. Even though they do not make the simplifying assumption, we should expect this error to be much larger. As mentioned earlier, the conditional dependence in such models changes in very specific ways (with conditional Kendall’s $\tau$ bounded by 1/3), and it wouldn’t be difficult to construct examples that deviate in bad ways. This comes on top of inherent symmetry properties of such models that are often violated in practice. For example, consider a hierarchical Archimedean model fit using the HAC package (Okhrin and Ristig,, 2014) with default settings to the simulated data from the previous section. While the $d_{\infty}$ distances from the true model are similar to the simplified vine, their KL divergence is substantially large:$\approx 0.89$ (HAC) vs. $\approx 0.67$ (vine) for the strong violation; $\approx 0.28$ (HAC) vs. $\approx 0.18$ (vine) for the moderate violation.
@@ 3.5 Simplified vines as statistical models
The first point highlights that we are rarely interested in the global approximation quality of $P_{\theta_{0}}$. If the goal is to estimate lower tail probabilities, we should not worry much about the approximation quality in the upper tail. If the goal is to predict $Y$ given covariates $\bm{X}$, we don’t care much about the marginal distributions of $\bm{X}$. If the goal is to derive simple explanations for the relationship between $Y$ and $\bm{X}$, a fully nonparametric model may be too complex to be useful. This urges statisticians to carefully consider what aspects of the model are important and which simplifications are justified to achieve their goal.
@@
The second point emphasizes that we cannot disentangle the approximation qualities of a model from our ability to estimate it. Generally speaking, more flexible models are harder to estimate—the ubiquitous bias-variance trade-off. In extreme cases, the model can be so large that consistent estimation is impossible. This is the case, for example, if $\Theta$ is the collection of all probability densities on $\mathds{R}^{d}$. Parametric families of densities on the other hand can be estimated at $O_{p}(n^{-1/2})$ rate under mild conditions. Merely assuming that the densities are twice continuously differentiable allows for estimation at $O_{p}(n^{-2/(d+4)})$ rate. This rate becomes extremely slow when $d$ is even moderately large. Nagler and Czado, (2016) showed that the simplifying assumption alleviates this. In particular, they construct nonparametric estimators of the partial vine copula that converge at $O_{p}(n^{-1/3})$ rate — irrespective of the dimension $d$. The simplifying assumption effectively makes estimation as easy as in a bivariate model. This comes at the cost of a potentially large approximation error, but it may be a good trade-off in practice. Nagler and Czado, (2016) show in simulations that, even on large samples, an estimated simplified model can be much closer to the true density than a fully nonparametric estimate, even if the simplifying assumption is violated severely. The gains in estimability simply outweigh the loss in approximation ability of the simplified model.
@@ 4 Conclusion
The simplifying assumption is a powerful tool for achieving computational and statistical efficiency while maintaining a high degree of flexibility. Of course, flexibility is to be understood in relative terms, compared to alternative models granting similar tractability. Simplified vines have proven immensely useful in numerous scientific fields. In many application areas, simplified vine copulas are among the most sophisticated and best fitting models known and accessible to field experts. Despite the simplifying assumption not always being true, these models are clearly a huge improvement over, for example, a multivariate Gaussian model. While this may sound like a low bar for people from the copula community, Gaussianity often forms the foundation of state-of-the-art methods in the application domain.
With part of the same reasoning, similar praise could be sung for generalized linear regression models. They are extremely useful in many fields but have their limitations. We know just about everything about these limitations. Which deviations cause which effect; how bad can a linear approximation be; how to test and correct for these deviations. The same cannot be said about the simplifying assumption. Practitioners tend to not worry about such an abstract assumption in an already complicated model. Copula experts may have a good intuition about what is happening, but intuition is not the same as scientific knowledge, which is currently scarce. Research that fills some of this gap is important and should be cherished.
@@
  • Test for the simplifying assumption if its validity is relevant (for example, with the pacotest R package, Kurz,, 2022). This is certainly the case when stakes are high or the interpretation of pair-copulas in higher trees is important.

### v3

@@ 1 Introduction
The simplifying assumption has become an oddly polarizing issue within our community, although the tension is not reflected in the published scientific record. Papers highlighting its limitations are dismissed in peer review for relying on ‘unrealistic toy examples.’ On the other side, methodological works on simplified models face rejections with the recommendation to abandon research involving the simplifying assumption altogether because it might be the ‘next formula that kills Wall Street’, referring to a famous article on the Gaussian copula (Salmon,, 2012). To my knowledge, neither of these positions has been voiced, let alone justified in the public scientific record. This state of affairs is quite unfortunate, as it hinders scientific progress.
The main purpose of this note is to bring the debate into the open scientific discourse and to offer my own perspective on the matter. At the very least, it should provide a good basis for a more constructive discourse going forward. The key points are as follows:
@@ 3.1.3 Extremal dependence
It appears that the simplifying assumption may sometimes be justified in the context of extremal dependence models. This should not be misinterpreted as a general endorsement of simplified vine copulas for estimating extreme event probabilities. Tail copulas are objects different from copulas and have to be fitted only on the most extreme subset of observations. A simplified vine model fit to all observations can still be off, because the estimated parameters may underrepresent the dependence strength in tail.
@@ 3.2 Empirical evidence for and against the simplifying assumption
The aforementioned works provide anecdotal evidence that whether and how severely the simplifying assumption is violated in practice depends on the data at hand. From a philosophical point of view, there isn’t much reason to believe it is exactly true—apart from the special case of conditional independence. Any consistent test will detect the smallest deviation if the sample size is large enough, and the above references all use at least several hundred observations both in real-data analyses and power studies.
The same can be said about any parametric model for the dependence and related goodness-of-fit tests. That does not mean that simplified or parametric models are inherently bad. What is important is that we understand the limitations of our models and the consequences of their violations.
@@ 3.3 Issues with interpreting simplified models
Figure 4: Contour plots of estimated pair-copulas in a simplified vine fit to non-simplified data (left: strong violation, right: moderate violation). While $X_{2}$ and $X_{3}$ are conditionally independent given $X_{1},X_{4}$, the partial vine copula induces spurious positive dependence (panels in the top row).
@@
  • Specify conditional copulas in the second tree with $c_{2,4|1}(\cdot\mid u)=c_{3,4|1}(\cdot\mid u)$ as a Gumbel copula with Kendall’s $\tau$ equal to $0.9(2u-1)$ (strong violation) or $0.4(2u-1)$ (moderate violation), rotating the copula by 90 degrees if $\tau<0$.
@@
We now simulate 10 0001000010\,00010 000 observations from the model and fit a nonparametric simplified vine copula using the rvinecopulib package (Nagler and Vatter,, 2024). The fitted model on the left shows strong positive dependence in the pair $(X_{2},X_{3})$ given $(X_{1},X_{4})$ (top row), even though they are conditionally independent. Also for the moderate violation (right plot), there is weak positive correlation with a Kendall’s $\tau$ of approximately 0.08, which passes usual statistical significance thresholds. Concluding that the variables are conditionally dependent would clearly be a mistake. For the moderate violation, the pair-copulas in the second tree further seem close to independence, because they average out positive and negative dependencies. Concluding that the variables are conditionally independent is also a mistake.
@@ 3.4 The approximation error of simplified models
The approximation error was analyzed formally in a seminal paper by Mroz et al., (2021). Their most surprising result is that the collection of simplified vine copulas is dense in the set of all copulas with respect to the uniform metric $d_{\infty}(C,D)=\sup_{\bm{u}\in[0,1]^{d}}|C(\bm{u})-D(\bm{u})|$. This means that the approximation error can be made arbitrarily small by choosing the right simplified model. We must not misinterpret this result as saying that the simplifying assumption is inconsequential in practice. The proof of the result is based on the fact that all empirical copulas can be represented as simplified vine copulas, with independence pair-copulas everywhere but for the first tree. The same argument would suggest that Markov tree distributions (e.g., Ansari and Ritter,, 2024) are dense in the set of all distributions. However, assuming universal conditional independence is clearly restrictive in a practical sense. And indeed, this all breaks down when more statistically meaningful metrics are considered: Mroz et al., (2021) show that simplified vine copulas are nowhere dense with respect to the topology induced by weak convergence of conditional distributions, the total variation metric, and the Kullback-Leibler divergence.
@@
It would be an equally big mistake to dismiss all models making the simplifying assumption based on these findings. The results are missing important context about alternative models. For example, I am not aware of similar results providing the approximation error of (hierarchical) Archimedean models, a popular alternative to vine copulas. Even though they do not make the simplifying assumption, we should expect this error to be much larger. As mentioned earlier, the conditional dependence in such models changes in very specific ways (with conditional Kendall’s $\tau$ bounded by 1/3), and it wouldn’t be difficult to construct examples that deviate in bad ways (for example, vine copulas containing conditional copulas with Kendall’s $\tau$ much larger than $1/3$ or much less than 00). This comes on top of inherent symmetry properties of such models that are often violated in practice. For example, consider a hierarchical Archimedean model fit using the HAC package (Okhrin and Ristig,, 2014) with default settings to the simulated data from the previous section. While the $d_{\infty}$ distances from the true model are similar to the simplified vine, their KL divergence is substantially large:$\approx 0.89$ (HAC) vs. $\approx 0.67$ (vine) for the strong violation; $\approx 0.28$ (HAC) vs. $\approx 0.18$ (vine) for the moderate violation. The purpose of this example is not to argue that vine copulas are better than Archimedean models, but to highlight that any model has its limitations and that the simplifying assumption is not the only one to be concerned about.
@@ 3.5 Simplified vines as statistical models
Point (i) highlights that we are rarely interested in the global approximation quality of $P_{\theta_{0}}$. If the goal is to estimate lower tail probabilities, we should not worry much about the approximation quality in the upper tail. If the goal is to predict $Y$ given covariates $\bm{X}$, we don’t care much about the marginal distributions of $\bm{X}$. If the goal is to derive simple explanations for the relationship between $Y$ and $\bm{X}$, a fully nonparametric model may be too complex to be useful. This urges statisticians to carefully consider what aspects of the model are important and which simplifications are justified to achieve their goal. A simple example where a deviation from the simplifying assumption may not be problematic is given in the following.
Example 3.1.
To give a simple example, where a deviation from the simplifying assumption is not problematic, consider the following regression model:
$Y=X_{1}+X_{2}\varepsilon,\quad\text{where}\quad(X_{1},X_{2},\varepsilon)\sim\mathcal{N}(0,I_{3}).$
The conditional distribution of $(Y,X_{1})$ given $X_{2}=x_{2}$ is a bivariate Gaussian with correlation $\rho(x_{2})=1/(1+x_{2}^{2})^{1/2}$, so the simplifying assumption is violated. However, the regression function is $\mathds{E}[Y\mid X_{1}=x_{1},X_{2}=x_{2}]=x_{1}$, which is the same as in the simplified model $(Y,X_{1},X_{2})\sim\mathcal{N}(0,\Sigma)$ with
$\Sigma=\begin{pmatrix}2&1&0\\ 1&1&0\\ 0&0&1\end{pmatrix}.$
In fact, both the marginal family for $Y$ and the copula are misspecified here. The situation changes if we are interested in conditional quantiles, where ignoring the simplifying assumption may lead to substantial errors.
Point (ii) emphasizes that we cannot disentangle the approximation qualities of a model from our ability to estimate it. Generally speaking, more flexible models are harder to estimate—the ubiquitous bias-variance trade-off. Specifically, suppose we are interested in some aspect $T$ of the true distribution $P_{0}$, such as extreme event probabilities or conditional quantiles. We can decompose the total error as follows:
$T(P_{\hat{\theta}_{n}})-T(P_{0})=\underbrace{T(P_{\widehat{\theta}_{n}})-T(P_{\theta_{0}})}_{\text{estimation error}}+\underbrace{T(P_{\theta_{0}})-T(P_{0})}_{\text{approximation error}},$
where $\widehat{\theta}_{n}$ is the estimated parameter based on $n$ observations. The estimation error usually decreases as the sample size $n$ grows, but the approximation error depends only on the model and the true distribution. Larger sample sizes thus allow us to use more flexible models, such as nonparametric or non-simplified ones.
@@
In extreme cases, the model can be so large that consistent estimation is impossible. This is the case, for example, if $\Theta$ is the collection of all probability densities on $\mathds{R}^{d}$. Parametric families of densities on the other hand can be estimated at $O_{p}(n^{-1/2})$ rate under mild conditions. Merely assuming that the densities are twice continuously differentiable allows for estimation at $O_{p}(n^{-2/(d+4)})$ rate. This rate becomes extremely slow when $d$ is even moderately large. Nagler and Czado, (2016) showed that the simplifying assumption alleviates this. In particular, they construct nonparametric estimators of the partial vine copula that converge at $O_{p}(n^{-1/3})$ rate — irrespective of the dimension $d$. The simplifying assumption effectively makes estimation as easy as in a bivariate model. This comes at the cost of a potentially large approximation error, but it may be a good trade-off in practice. Nagler and Czado, (2016) show in simulations that, even on large samples, an estimated simplified model can be much closer to the true density than a fully nonparametric estimate, even if the simplifying assumption is violated severely. The gains in estimability may simply outweigh the loss in approximation ability of the simplified model.
@@ 4 Conclusion
The simplifying assumption is a powerful tool for achieving computational and statistical efficiency while maintaining a high degree of flexibility. Of course, flexibility is to be understood in relative terms, compared to alternative models granting similar tractability. Simplified vines have proven immensely useful in numerous scientific fields. In many application areas, simplified vine copulas are among the most sophisticated and best fitting models known and accessible to field experts. Despite the simplifying assumption not always being true, these models are clearly a huge improvement over, for example, a multivariate Gaussian model (which also satisfies the simplifying assumption). While this may sound like a low bar for people from the copula community, Gaussianity often forms the foundation of state-of-the-art methods in the application domain.
With part of the same reasoning, similar praise could be sung for generalized linear regression models. They are extremely useful in many fields but have their limitations. We know just about everything about these limitations. Which deviations cause which effect; how bad can a linear approximation be; how to test and correct for these deviations. The same cannot be said about the simplifying assumption. Practitioners tend to not worry about such an abstract assumption in an already complicated model. Copula experts may have a good intuition about what is happening, but intuition is not the same as scientific knowledge, which is currently scarce. A non-exhaustive list of important questions that remain to be answered is:
  • How bad are violations of the simplifying assumption if focusing on parametric models?
  • How do we appropriately measure/estimate the approximation error of a specific simplified vine copula model to the true copula?
  • How can we mitigate the effects of violations of the simplifying assumption in a computationally and statistically efficient way?
  • Which downstream tasks are insensitive to violations of the simplifying assumption, which are not?
Research that fills some of this gap is important and should be cherished.
@@
  • Test for the simplifying assumption if its validity is relevant. This is certainly the case when stakes are high or the interpretation of pair-copulas in higher trees is important. Several R packages are available for testing the assumption, visual diagnostics, or non-simplified vine models; see Kurz, (2022), Acar and Lysy, (2024), Nagler and Vatter, (2020), Derumigny, (2024).